# CoSy: Evaluating Textual Explanations of Neurons

### v1

@@ A.6 Model Stability
For our analysis, we utilize ResNet18 and focus on its output neurons, as the ground-truth labels associated with these neurons are known. We randomly select six classes $s$ from the ImageNet validation dataset [russakovsky2015imagenet] and examine the corresponding $f_{i}$ class output neurons using CoSy. Here, we exclude the $s$ class from $\mathbb{A}_{0}$ and let $\mathbb{A}_{1}$ represent the $s$ class. To ensure robustness, we initialize the text-to-image model across a random set of 10 seeds. Our analysis involves calculating the first (mean) and second moment (STD) using $\Psi_{\text{AUC}}$, as well as evaluating the intraclass image similarity (refer to Section A.5) within each synthetic ground truth class.
@@
Table 6: Model Stability. A comparison of various model initializations across 10 random seeds using SDXL.
Concept AUC ($\uparrow$) CS ($\uparrow$) ED ($\downarrow$)
bulbul 0.9996$\pm$0.0002 0.91$\pm$0.03 3.99$\pm$0.66
china cabinet 0.9999$\pm$0.0001 0.89$\pm$0.04 5.00$\pm$0.90
leatherback turtle 0.9994$\pm$0.0001 0.91$\pm$0.04 4.65$\pm$0.87
beer bottle 0.9919$\pm$0.0038 0.80$\pm$0.08 6.79$\pm$1.41
half track 0.9998$\pm$0.0000 0.88$\pm$0.04 5.12$\pm$0.91
hard disc 1.0000$\pm$0.0001 0.90$\pm$0.05 4.64$\pm$1.17
Overall Mean 0.9984$\pm$0.0007 0.88$\pm$0.02 5.03$\pm$0.26
A.7 Prompt and Text-to-Image Model Comparison
@@
Figure 8: Example images for “coffee mug” generated by the text-to-image models SDXL and SC across various prompts. (1) and (3) present examples of synthetic images with relatively low intraclass similarity and relatively high natural-to-synthetic similarity scores. (2) shows examples of synthetic images with the lowest similarity to natural images. (4) illustrates examples of synthetic images with the highest similarity to other synthetic images within the same class. (5) showcases examples of synthetic images with the highest similarity to natural images. (6) displays examples of natural images from the ImageNet validation dataset [russakovsky2015imagenet] belonging to the class “coffee mug”.

### v2

@@ A.6 Model Stability
For our analysis, we utilize ResNet18 and focus on its output neurons, as the ground-truth labels associated with these neurons are known. We randomly select six classes $s$ from the ImageNet validation dataset [36] and examine the corresponding $f_{i}$ class output neurons using CoSy. Here, we exclude the $s$ class from $\mathbb{A}_{0}$ and let $\mathbb{A}_{1}$ represent the $s$ class. To ensure robustness, we initialize the text-to-image model across a random set of 10 seeds. Our analysis involves calculating the first (mean) and second moment (STD) using $\Psi_{\text{AUC}}$, as well as evaluating the intraclass image similarity (refer to Section A.5) within each synthetic ground truth class.
@@
Table 8: Model Stability. A comparison of various model initializations across 10 random seeds using SDXL. The results represent the average scores with standard deviations for each class, calculated across 10 seeds.
Concept AUC ($\uparrow$) Similarity Measure
CS ($\uparrow$) ED ($\downarrow$) LPIPS ($\downarrow$)
bulbul 0.9996$\pm$0.0002 0.91$\pm$0.03 3.99$\pm$0.66 0.65$\pm$0.06
china cabinet 0.9999$\pm$0.0001 0.89$\pm$0.04 5.00$\pm$0.90 0.58$\pm$0.04
leatherback turtle 0.9994$\pm$0.0001 0.91$\pm$0.04 4.65$\pm$0.87 0.63$\pm$0.04
beer bottle 0.9919$\pm$0.0038 0.80$\pm$0.08 6.79$\pm$1.41 0.69$\pm$0.08
half track 0.9998$\pm$0.0000 0.88$\pm$0.04 5.12$\pm$0.91 0.61$\pm$0.04
hard disc 1.0000$\pm$0.0001 0.90$\pm$0.05 4.64$\pm$1.17 0.60$\pm$0.05
Overall Mean 0.9984$\pm$0.0007 0.88$\pm$0.02 5.03$\pm$0.26 0.63$\pm$0.05
A.7 Compute Resources
A.8 Additional Results for Method Comparison across ResNet18 Layers
Given that the original implementation of FALCON only provides results for their defined “explainable neurons” (see Appendix A.1.4), we included additional results comparing all methods based on this subset of neurons. Specifically, there are 7 explainable neurons in layer 2, 5 in layer 3, and 15 in layer 4. Figure 8 presents these results.
Figure 8: A comparison of explanation methods in ResNet18 shows that INVERT and CLIP-Dissect maintain high MAD scores across all layers, while MILAN and FALCON have lower scores. Overall, performance declines in the lower layers for all methods.
A.9 Qualitative Examples of Lower-Level Concepts
The performance of generative models may present challenges when dealing with abstract concepts. To assess this concern and provide a basis for further discussion, we have included an additional example of lower-layer concepts in Figure 9.
Figure 9: Examples of various explanations from four different methods applied to the low-layer neurons in layer 2 of the ResNet18 model are provided. From the illustrated Feature Visualizations, we can observe that these neurons detect low-level abstractions. However, the methods studied generally fail to provide low-level explanations, instead attributing more complex explanations.
A.10 Concept Broadness
While CoSy focuses on measuring the explanation quality, another open question is how broad or abstract are the concepts provided as textual explanations. This question of how specific or general an individual neuron is described by the explanation, might be relevant to different interpretability applications. For example, research fields where the user aims to deploy the same network for multiple tasks with varying image domains. In this case, describing a neuron’s more general concept such as “a round object” might be more informative than a more (domain-)specific concept such as “a tennis ball” for the network assessment. To provide insight into the broadness of concepts, we assessed whether the similarity between images generated based on the same concept changes from more general to more specific concepts.
In our experiment, we define the broadness of a concept based on the number of hypernyms in the WordNet hierarchy [54]. The more specific a concept the larger the number of hypernyms. We choose two ImageNet classes (“ladybug”, “pug”) and generate $50$ images for each concept as well as each hypernym of both concepts (with the most general concept being “entity”). Then, we measure the Cosine Similarity (CS) of all images generated based on the same concept. The box plot of the CS across both concepts and all hypernyms, in Figure 10 indicates that we do not find a correlation. Thus, we hypothesize that the chosen temperature of the diffusion model has a stronger effect on image similarity than the broadness of the prompt used for image generation.
Figure 10: The figure demonstrates the independence of the concept broadness measured by the number of hypernyms as defined in WordNet [54] to the inter-image similarity of corresponding generated images.
A.11 Prompt and Text-to-Image Model Comparison
@@
Figure 11: Example images for “coffee mug” generated by the text-to-image models SDXL and SC across various prompts. (1) and (3) present examples of synthetic images with relatively low intraclass similarity and relatively high natural-to-synthetic similarity scores. (2) shows examples of synthetic images with the lowest similarity to natural images. (4) illustrates examples of synthetic images with the highest similarity to other synthetic images within the same class. (5) showcases examples of synthetic images with the highest similarity to natural images. (6) displays examples of natural images from the ImageNet validation dataset [36] belonging to the class “coffee mug”.